Descibe pattern. Then write the next three numbers.
1) 640,320,160,80...
2) 4,12,36,108...
3) 1,1,2,3,5,8...
4) 2,11,38,119...

Answer by Edwin McCravy(8999)

Descibe pattern. Then write the next three numbers.
1) 640,320,160,80...

Start with 640, 
Take half of 640, get 320
Take half of 320, get 160
Take half of 160, get 80
So,
Take half of 80, get 40
Take half of 40, get 20
Take half of 20, get 10


2) 4,12,36,108...

Start with 4, 
Multiply 4 by 3, get 12
Multiply 12 by 3, get 36
Multiply 35 by 3, get 108
So,
Multiply 108 by 3, get 324
Multiply 324 by 3, get 972
Multiply 972 by 3, get 2916


3) 1,1,2,3,5,8...

Start with the first two 1,1
To get the 3rd term, add the 1st and 2nd, get 1+1=2
To get the 4th term, add the 2nd and 3rd, get 1+2=3
To get the 5th term, add the 3rd and 4th, get 2+3=5
So
To get the 6th term, add the 4th and 5th, get 3+5=8
To get the 7th term, add the 5th and 6th, get 5+8=13
To get the 8th term, add the 6th and 7th, get 8+13=21


4) 2,11,38,119...

To get the 1st term, 
1. multiply 1 by 6, get 1×6 = 6
2. Subtract that from 27, get 27-6 = 21,
3. Multiply that by 1, get 21×1 = 21
4. Subtract that from 48, get 48-21 = 27
5. Multiply that by 1, get 27×1 = 27
6. Subtract 25 from that, get 27-25 = 2, the 1st term.


To get the 2nd term, 
1. multiply 2 by 6, get 2×6 = 12
2. Subtract that from 27, get 27-12 = 15,
3. Multiply that by 2, get 15×2 = 30
4. Subtract that from 48, get 48-30 = 18
5. Multiply that by 2, get 18×2 = 36
6. Subtract 25 from that, get 36-25 = 11, the 2nd term

To get the 3rd term, 
1. multiply 3 by 6, get 3×6 = 18
2. Subtract that from 27, get 27-18 = 9,
3. Multiply that by 3, get 9×3 = 27
4. Subtract that from 48, get 48-27 = 21
5. Multiply that by 3, get 21×3 = 63
6. Subtract 25 from that, get 63-25 = 38, the 3rd term

To get the 4th term, 
1. multiply 4 by 6, get 4×6 = 24
2. Subtract that from 27, get 27-24 = 3,
3. Multiply that by , get 3×4 = 12
4. Subtract that from 48, get 48-12 = 36
5. Multiply that by 4, get 36×4 = 144
6. Subtract 25 from that, get 144-25 = 119, the 4th term


So,

To get the 5th term, 
1. multiply 5 by 6, get 5×6 = 30
2. Subtract that from 27, get 27-30 = -3,
3. Multiply that by 5, get -3×5 = -15
4. Subtract that from 48, get 48-(-15) = 48+15 = 63
5. Multiply that by 5, get 63×5 = 315
6. Subtract 25 from that, get 315-25 = 290, the 5th term

To get the 6th term, 
1. multiply 6 by 6, get 6×6 = 36
2. Subtract that from 27, get 27-36 = -9,
3. Multiply that by 6, get -9×6 = -54
4. Subtract that from 48, get 48-(-54) = 48+54 = 102
5. Multiply that by 6, get 102×6 = 612
6. Subtract 25 from that, get 612-25 = 587, the 6th term

To get the 7th term, 
1. multiply 7 by 6, get 7×6 = 42
2. Subtract that from 27, get 27-42 = -15,
3. Multiply that by 7, get -15×7 = -105
4. Subtract that from 48, get 48-(-105) = 48+105 = 153
5. Multiply that by 7, get 153×7 = 1071
6. Subtract 25 from that, get 1071-25 = 1046, the 7th term
 
Edwin